# Towards Contextual Sensitive Data Detection

### v1

@@ Abstract
The emergence of open data portals necessitates more attention to protecting sensitive data before datasets get published and exchanged. While an abundance of methods for suppressing sensitive data exist, the conceptualization of sensitive data and methods to detect it, focus particularly on personal data that, if disclosed, may be harmful or violate privacy. We observe the need for refining and broadening our definitions of sensitive data, and argue that the sensitivity of data depends on its context. Based on this definition, we introduce two mechanisms for contextual sensitive data detection that consider the broader context of a dataset at hand. First, we introduce type contextualization, which first detects the semantic type of particular data values, then considers the overall context of the data values within the dataset or document. Second, we introduce domain contextualization which determines sensitivity of a given dataset in the broader context based on the retrieval of relevant rules from documents that specify data sensitivity (e.g., data topic and geographic origin). Experiments with these mechanisms, assisted by large language models (LLMs), confirm that: 1) type-contextualization significantly reduces the number of false positives for type-based sensitive data detection and reaches a recall of 94% compared to 63% with commercial tools, and 2) domain-contextualization leveraging sensitivity rule retrieval is effective for context-grounded sensitive data detection in non-standard data domains such as humanitarian datasets. Evaluation with humanitarian data experts also reveals that context-grounded LLM explanations provide useful guidance in manual data auditing processes, improving consistency. We open-source mechanisms and annotated datasets for contextual sensitive data detection at https://github.com/trl-lab/sensitive-data-detection.
@@ 1 Introduction
Data sharing underpins reuse across sciences, enterprises, and governments through public and private portals [borgman2025data, worth2024datatransparency, zhang2024datamarketssurvey, wilkinson2016fair, brickley2019googledataset]. With the rise of open data sharing portals, comes the need to protect sensitive data from undesired and harmful use. Sensitive data includes direct personal identifiers (e.g., names, emails [gdpr]) as well as indirect or contextual information that becomes a potential risk under certain conditions [unocha2025]. For example, facility locations or military logs may be sensitive depending on geopolitics or timing. This illustrates the complexity of sensitive data detection and the need for nuanced, context-dependent methods. Beyond operational and ethical concerns of publishing sensitive data, Large Language Models (LLMs) increase the risks of leakage: trained on web data, they may memorize and reproduce sensitive information [subramani2023piihuggingface, worth2024datatransparency, carlini2021extracting, lukas2023analyzing]. Studies show sensitive data is often underspecified in documentation on open data portals like HuggingFace [yang2024navigatingdatasetdocumentationsai, croissant2024benjelloun, worth2024datatransparency], illustrating the importance of accurate methods and tools for (semi-)automated sensitive data protection, particularly in open data portals.
@@
Most work on sensitive data detection methods, however, focuses on detecting person-level data, particularly personal identifiable information (PII) [KUZINA2023119924, subramani2023piihuggingface, croissant2024benjelloun]. For unstructured data, PII detection relies on named entity recognition [subramani2023piihuggingface]; similar approaches exist for tabular data [raman2001potter, Hulsebos2019Sherlock], which is the focus in this paper. Yet sensitivity extends far beyond explicit PII. kober2023whatissensitive propose a taxonomy of always-, combination-, context-, and value-sensitive data. For instance, an address may be harmless in the context of a public organizational website but expose a private address in another document, causing naïve type-based detection yield false positives. Likewise, hospital geo-coordinates may be harmless in one region but sensitive in conflict zones [unocha2025], calling for contextualization beyond detecting personal-sensitive data. These limitations call for a more nuanced perspective and framework for assessing sensitive data. LLMs offer new potential for sensitivity detection by reasoning over structured data. Studies show they can interpret tables with carefully designed prompts [fang2024largelanguagemodelsllmstabular], achieving strong results in tasks such as text-to-SQL and table retrieval [fang2024largelanguagemodelsllmstabular, ji2025target]. These capabilities suggest that LLMs can leverage schema, values, and external knowledge to interpret tabular dataset semantics [Hulsebos2019Sherlock].
@@
In this paper, we first revisit the concept of sensitive data, arguing for a more contextual perspective. We introduce two new LLM-assisted mechanisms for contextual sensitive data detection. First, we introduce the detect-then-reflect mechanism for type contextualization of data values within the entire document or tabular dataset to assess their sensitivity. Experiments with real-world datasets containing personal identifiable information show that this mechanism yields a strong improvement in precision while maintaining high recall. Second, the retrieve-then-detect mechanism for domain contextualization which grounds the assessment in domain-specific sensitivity specifications. Evaluations illustrate the potential of this mechanism, achieving high recall and improved consistency for detecting non-personal sensitive data. Importantly, the context-grounded domain-specific explanations were deemed highly useful by experts in humanitarian data assessment in a case-study.
2 Background
Conceptualization of Sensitive Data
@@ 2 Background
The notion of “sensitive data” is multifaceted, spanning legal, theoretical, and domain-specific views. Legal frameworks often equate sensitivity with personal identifiable information. The GDPR defines personal data broadly as “any information relating to an identified or identifiable natural person,” specifying sensitive categories like health, ethnicity, or politics requiring extra protection [gdpr]. Similarly, HIPAA defines identifiers that make health information “Protected Health Information (PHI)” when linked to individuals [hipaa]. These protocols provide a baseline: data that directly or indirectly identifies a person is sensitive. Yet static, type-level definitions are insufficient for modern governance [Quinn_Malgieri_2021], as they emphasize direct identifiers while overlooking non-personal but risky data such as geolocation data, community-level attributes, or institutional records. Sensitivity is often contextual: seemingly harmless data can become sensitive when combined, repurposed, or shared across contexts [malkin2022contextualintegrity].
Theoretical perspectives deepen this nuance. Nissenbaum’s Contextual Integrity frames sensitivity as the appropriateness of information flows within social norms [nissenbaum2004privacy]. Violations occur when flows breach expectations, not from the data type itself. Building on this, researchers advocate context-grounded definitions [malkin2022contextualintegrity, shvartzshnaider2019going]. For instance, shopping data may be ordinary in retail but sensitive in behavioral profiling. Recent taxonomies further refine this view. kober2023whatissensitive distinguish: (1) always-sensitive data (e.g., direct identifiers); (2) combination-sensitive (sensitive when linked together); (3) context-sensitive (sensitivity by use); and (4) value-sensitive (sensitivity from specific values). This shows that not all personal is equally sensitive, and not all sensitive data is on individuals.
In short, conceptualizations have shifted from static identifiers to context-dependent models, where frameworks like Contextual Integrity and multi-layer taxonomies emphasize context, use, and value. This motivates our focus: effective data protection must ensure coverage beyond personal-level information toward context-aware, domain-specific approaches.
Methods for Detecting Sensitive Data
Most automated sensitive data detection methods concentrate on detecting sensitive data of individuals. Widely adopted industry tools like Microsoft Presidio [presidio] and Google Cloud DLP [google_DLP] are largely rule-based, relying on regular expressions, keyword lists, and simple machine learning heuristics to recognize well-defined patterns like email addresses or phone numbers. Research has advanced towards more semantic understanding. Hulsebos2019Sherlock introduced Sherlock, a deep learning model that classifies types of table columns, including numerous PII, by learning from column-level representations. KUZINA2023119924 develop this further with CASSED, a BERT-based model that integrates column headers, data types, and sample values to classify columns into sensitive categories. These approaches outperform regular expressions and dictionary-based baselines, especially on noisy or inconsistently formatted data. Recently, Large Language Models (LLMs) have emerged as powerful tools for sensitive data detection, as particularly shown for PII detection outperforming both rule-based and traditional machine learning methods in accuracy and adaptability[shen2025enhancingdeidentificationpersonallyidentifiable, Yang2023Exploring]. However, a key limitation remains: LLMs excel at identifying known PII types but struggle with assessing contextual sensitivity, such as domain-specific sensitive types, unless explicitly provided with the relevant operational context.
We identify a significant gap in detecting sensitive data: methods focus on detecting personal sensitive data, overlooking non-personal data that is sensitive due to its context. Our work bridges this gap by developing a framework that addresses both personal and non-personal sensitive data detection through contextualization.
3 Revisiting the Concept of Sensitive Data
The main conceptual, legal, and methodological focus regarding sensitive data is on personal- and individual level data, including personal identifiable information (PII) which refers to data that can identify an individual [gdpr]. Regulatory data protocols, such as GDPR [gdpr] and HIPAA [hipaa], specify standards for data sensitivity, hence provide concrete guidelines for data protection measures. However, we argue for the need of a more contextual view of sensitive data (Def. 1) for developing new detection methods.
Definition 1(Contextual Sensitive Data)
@@ 3 Revisiting the Concept of Sensitive Data
Advancing this definition, we posit two contextual aspects that matter in order to assess sensitivity of data: type contextualization and domain contextualization.
Following the existence of generic standards and protocols, sensitive data detection methods often focus on detecting data types reflecting personal sensitive data (e.g., names, addresses, phone numbers, birth dates). But the appearance of a name or address does not immediately make a data value sensitive. For example, an address may reflect the address of a person or organization, each having significantly different consequences, hence different sensitivity levels. We term this type contextualization: whether a data value corresponding to a sensitivity type of interest (e.g. name) truly is sensitive requires considering the context in which it appears in the entire dataset (e.g., the entire document or table).
Besides a more fine-grained contextual notion of a data value its sensitivity, we also argue for a more coarse-grained contextual perspective. Nissenbaum introduced the theory of Contextual Integrity [nissenbaum2004privacy], which states that privacy risks arise when data is used outside its intended setting, particularly addressing personal privacy. This contextual perspective extends beyond personal privacy, in domains such as humanitarian aid, national security, and company-sensitive context. We pose that data protection should account for risks beyond identity exposure, for example, harming vulnerable populations or organizations. We refer to this broader perspective on data sensitivity as domain contextualization: sensitivity emerges from the dataset’s nature and potential misuse.
4 Toward Contextual Sensitive Data Detection
Protecting sensitive data generally follows a three-step procedure: ➀ specifying sensitive data, ➁ detecting sensitive data, ➂ remedying data sensitivity (Figure 1). In this work, we focus on the second step, which has received little attention in research. Existing methods for detecting sensitive data suffer from two key limitations to address the contextual nature of sensitive data:
@@ 4 Toward Contextual Sensitive Data Detection
    Insufficient specificity: sensitive data detection methods treat all values of types of interest as equally sensitive, while the subject of the value that implies its sensitivity. The lack of specificity leads to low precision hence high false positive rates.
@@
    Insufficient coverage: sensitive data detection methods typically focus on personal-level data, particularly PII, and do not generalize to domain-specific sensitive data. This lack of coverage leads to low recall of methods, hence leaving sensitive data unprotected.
Figure 1: Data protection process, starting with the specification of sensitivity, detection sensitive data in inputs accordingly, followed by remedying sensitive data by suppression methods.
@@
The contextualization framework introduced in Section 3 addresses these limitations. Type contextualization would improve specificity as potential sensitive values are only considered sensitive based on the broader context of the value. In turn, domain contextualization ensures higher coverage of sensitive data assessment by grounding the assessment in domain-specific context. In this paper, we introduce two concrete mechanisms to address these contextual aspects of sensitive data: the detect-then-reflect mechanism for type contextualization and the retrieve-then-detect mechanism for domain contextualization. In what follows, we specifically focus on tabular data but these mechanisms extend to other modalities, such as free-form text documents.
@@
(a) The detect-then-reflect mechanism for detecting sensitive table columns using type contextualization to increase the precision. This mechanism entails: ➀ detecting potentially sensitive columns based on their type, ➁ having an LLM reflect on the actual sensitivity of the column within the context of the entire table.
@@
(b) The retrieve-then-detect mechanism for detecting sensitive data that requires domain contextualization. This mechanism entails: ➀ retrieving relevant contextual information such as rule-based heuristics, ➁ using an LLM to detect sensitive data using its parametric knowledge and the retrieved expert heuristics.
Figure 2: Mechanisms for detecting sensitive data in a contextual manner.
4.1 Type contextualization: detect-then-reflect
Existing methods typically classify sensitive data based on signals indicating a type of interest, e.g. names, addresses, or phone numbers. If a value matches this type, it is flagged as sensitive. We introduce a two-step pipeline for detecting type-informed sensitive data based on a detect-then-reflect procedure (Figure 2(a)). In the initial detection step, potentially sensitive columns are identified by first classifying columns corresponding to types of interest that are specified upfront, e.g. email, phone number, based on column names and data values. This step prioritizes recall in order to avoid missing any columns that may carry data of the type that can potentially be sensitive. In the second step, reflection, the model assesses the columns in the context of the full table. We prompt an LLM with the table, including the column’s detected type. The LLM determines whether the column is truly sensitive, taking into account inter-column relationships and data across the entire table. This step is crucial for mitigating false positives from the detection stage by incorporating inter-column relationships and table semantics.
4.2 Domain contextualization: retrieve-then-detect
While type-based data sensitivity can often be inferred from structural context, the data sensitivity may depend on domain-specific and geopolitical factors that lie beyond the data itself. For such cases, requiring domain contextualization, we introduce the retrieve-then-detect mechanism (Figure 2(b)). This mechanism considers external contextual information, such as rule-based heuristics specified in data governance documents, in the classification process. Relevant documents are retrieved based on, for example, the origin of a dataset and included in the prompt alongside the table and column under analysis. The LLM leverages both its internal parametric knowledge and the retrieved external data sensitivity instructions to assess whether a column contains data that should be considered sensitive in the given operational context. Unlike type-informed detection, where there is a predefined list of entities (e.g. email, phone number) as relevant for for example PII detection, domain-contextual sensitive data has no fixed ontology. This mechanism establishes domain-specific column sensitivity based on the broader table context and retrieved contextual knowledge.

### v2

@@ Abstract
The emergence of open data portals necessitates more attention to protecting sensitive data before datasets get published and exchanged. To do so effectively, we observe the need to refine and broaden our definitions of sensitive data, and argue that the sensitivity of data depends on its context. Following this definition, we introduce a contextual data sensitivity framework building on two core concepts: 1) type contextualization, which considers the type of the data values at hand within the overall context of the dataset or document to assess their true sensitivity, and 2) domain contextualization, which assesses the sensitivity of data values informed by domain-specific information external to the dataset, such as geographic origin of a dataset. Experiments instrumented with language models confirm that: 1) type-contextualization significantly reduces the number of false positives for type-based sensitive data detection and reaches a recall of 94% compared to 63% with commercial tools, and 2) domain-contextualization leveraging sensitivity rule retrieval effectively grounds sensitive data detection in relevant context in non-standard data domains. A case study with humanitarian data experts also illustrates that context-grounded explanations provide useful guidance in manual data auditing processes. We open-source the implementation of the mechanisms and annotated datasets at https://github.com/trl-lab/sensitive-data-detection.
Towards Contextual Sensitive Data Detection
Liang Telkamp1,2 and Madelon Hulsebos2thanks: Correspondence to madelon.hulsebos@cwi.nl 1University of Amsterdam, 2CWI
@@ 1 Introduction
Data sharing underpins data reuse across sciences, enterprises, and governments through public and private portals (Borgman and Groth, 2025; Worth et al., 2024; Zhang et al., 2024; Wilkinson et al., 2016; Brickley et al., 2019). With the rise of open data sharing portals, comes the need to protect sensitive data from undesired and harmful use. Sensitive data includes direct personal identifiers (e.g., names, emails (European Parliament and Council of the European Union, 2016)) as well as indirect or contextual information that becomes a potential risk under certain conditions (OCHA, 2025). For example, facility locations or military logs may be sensitive depending on geopolitics or timing. This illustrates the complexity of sensitive data detection and the need for nuanced, context-dependent methods. Beyond operational and ethical concerns of publishing sensitive data, Large Language Models (LLMs) increase the risks of leakage: trained on web data, they may memorize and reproduce sensitive information (Subramani et al., 2023; Worth et al., 2024; Carlini et al., 2021; Lukas et al., 2023). Studies show sensitive data is often underspecified in documentation on open data portals like HuggingFace (Yang et al., 2024; Akhtar et al., 2024; Worth et al., 2024), illustrating the importance of accurate methods and tools for (semi-)automated sensitive data protection, particularly on open portals.
@@
Most work on sensitive data detection methods, however, focuses on detecting person-level data, particularly personal identifiable information (PII) (Kužina et al., 2023; Subramani et al., 2023; Akhtar et al., 2024). For unstructured data, PII detection relies on named entity recognition (Subramani et al., 2023); similar approaches exist for tabular data (Raman, 2001; Hulsebos et al., 2019), which is the focus in this paper. Yet sensitivity extends far beyond explicit PII. Kober et al. (2023) propose a taxonomy of always-, combination-, context-, and value-sensitive data. For instance, an address may be harmless in the context of a public organizational website but expose a private address in another document, causing naïve type-based detection yield false positives. Likewise, hospital geo-coordinates may be harmless in one region but sensitive in conflict zones (OCHA, 2025), calling for contextualization beyond detecting personal-sensitive data. These limitations call for a more nuanced perspective and framework for assessing sensitive data.
@@
In this paper, we revisit the concept of sensitive data detection, arguing for a more contextual perspective. We particularly focus on two aspects of contextual sensitive data detection. First, type contextualization which takes a more nuanced view beyond the type of values (e.g. addresses) by taking the broader context of the data into account. We introduce the technical “detect-then-reflect” mechanism which first detects the type of the values at hand, then reflects on their actual sensitivity based on contextual signals within the document or dataset. Experiments with real-world datasets containing personal identifiable information (PII) show that this mechanism, instrumented by language models, yields a strong improvement in precision while maintaining high recall. Second, we posit domain contextualization which addresses sensitivity that depends on external domain-specific factors, beyond merely static types such as PII. We introduce the “retrieve-then-detect” mechanism which first retrieves relevant domain-specific information, such as applicable data policies, to then ground the data sensitivity detection in using a language model. Evaluations illustrate the potential of this mechanism, achieving high recall and improved consistency for detecting non-personal sensitive data. Importantly, the context-grounded domain-specific explanations are deemed highly useful by experts in humanitarian data assessment in a case-study.
2 Revisiting Sensitive Data Detection
@@
The notion of “sensitive data” is multifaceted, spanning legal, theoretical, and domain-specific views. Legal frameworks often equate sensitivity with personal identifiable information. The GDPR defines personal data broadly as “any information relating to an identified or identifiable natural person,” specifying sensitive categories like health, ethnicity, or politics requiring extra protection (European Parliament and Council of the European Union, 2016). Similarly, HIPAA defines identifiers that make health information “Protected Health Information (PHI)” when linked to individuals (Centers for Medicare & Medicaid Services, 1996). Yet, sensitivity is often contextual: seemingly harmless data can become sensitive when combined, repurposed, or shared across contexts (Malkin, 2022).Theoretical perspectives deepen this nuance. Nissenbaum’s Contextual Integrity frames sensitivity as the appropriateness of information flows within social norms (Nissenbaum, 2004). Building on this, researchers advocate context-grounded definitions (Malkin, 2022).Kober et al. (2023) distinguish: (1) always-sensitive data (e.g., personal identifiers); (2) sensitivity by linking data; (3) sensitivity by use; and (4) sensitivity from specific values. This echos that not all personal is equally sensitive, and not all sensitive data is on individuals. To detect sensitive data, of any kind, our framework builds on this conceptual notion of data sensitivity (Def. 1): data protection methods must ensure coverage beyond personal- and static sensitivity definitions.
Definition 1 (Contextual Sensitive Data).
@@
Informed by this contextual definition of sensitive data, we formalize the problem of contextual sensitive data detection as follows:
Problem Statement 1 (Contextual Sensitive Data Detection).
Given an unordered or ordered sample of values $V={v_{1},v_{2},…,v_{N}}$ from document D, where each value $v_{i}$ represents a value, the set of internal context factors $C_{I}$ within $\mathcal{D}$ (e.g., surrounding values, structure), and external context factors $C_{E}$ outside $\mathcal{D}$ (e.g., specified rules, time/location), the objective of contextual sensitive data detection is to estimate the probability $P(V=sensitive|C_{E},C_{I})\in[0,1]$, and, based on this probability, classify the sensitivity of V into one of the predefined sensitivity levels $s_{1},s_{2},…,s_{M}$}, where each $s_{i}$ corresponds to a distinct sensitivity level (e.g., "sensitive", “medium sensitive”, "non-sensitive", etc.).
3 Contextual Sensitive Data Detection
Protecting sensitive data generally follows a three-step procedure: ➀ specifying sensitive data, ➁ detecting sensitive data, ➂ remedying data sensitivity. In this work, we focus on the second step, which has received little attention in research. Existing methods for detecting sensitive data suffer from two key limitations to address the contextual nature of sensitive data:
@@ 3 Contextual Sensitive Data Detection
    Insufficient specificity: sensitive data detection methods treat all values of types of interest as equally sensitive, while the subject of the value implies its sensitivity. This indistinctness leads to low precision hence high false positive rates.
@@
    Insufficient coverage: sensitive data detection methods typically focus on personal-level data, particularly PII, and do not generalize to domain-specific sensitive data. The lack of coverage leads to low recall of methods, hence leaving sensitive data unprotected.
@@
To address these challenges in contextual sensitive data detection, we posit two contextual aspects that matter in order to assess the contextual sensitivity of data: type contextualization and domain contextualization. We introduce two mechanisms to consider these contextual aspects of sensitive data: the detect-then-reflect mechanism for type contextualization and the retrieve-then-detect mechanism for domain contextualization. In what follows, we specifically focus on tabular data but these mechanisms extend to other modalities, such as free-form text documents.
3.1 Type Contextualization
Following existing standards and protocols European Parliament and Council of the European Union (2016); Centers for Medicare & Medicaid Services (1996), sensitive data detection methods often focus on detecting data types reflecting personal sensitive data (e.g., names, addresses) Google (2018). But the appearance of a name or address does not immediately make a data value sensitive. For example, an address may reflect the address of a person or organization, each having significantly different consequences, hence different sensitivity levels. We term this type contextualization: whether a data value corresponding to a sensitivity type of interest (e.g. name) truly is sensitive requires considering the context within the document or dataset in which it appears. Type contextualization improves specificity as potential sensitive values are only considered sensitive based on the broader context of the value.
The detect-then-reflect mechanism
Existing methods classify sensitive data based on signals indicating a type of interest, e.g. names, addresses, or phone numbers. If a value matches this type, it is flagged as sensitive. Type contextualization employs a more nuanced assessment, for which we introduce the two-stage detect-then-reflect mechanism for detecting type-informed sensitive data. Algorithm 1 illustrates this staged procedure for binary sensitivity classification.
Algorithm 1 Type Contextualization
1:Values $V_{i}=\{v_{1},\dots,v_{N}\}$ from document $D$
2:Internal context $C_{I}$, sensitive types $T_{\text{sens}}$, threshold $t$
[…]
4:if $t_{i}\in T_{\text{sens}}$ then
5:  $p_{\text{sens}}\leftarrow\textsc{Estimate}$ $P(V_{i}=\text{sensitive}\mid t_{i},C_{I})$
6:  $s_{i}\leftarrow(p_{\text{sens}}>t)$
7:else
8:  $s_{i}\leftarrow$ false
9:end if
In the initial detection step, the potential sensitivity of a set of values is estimated by first classifying the values ($V_{i}$) corresponding to sensitive types of interest specified upfront ($T_{\text{sens}}$). This step prioritizes recall by capturing all values that may carry data of the sensitive type. The purpose of the reflection step is to estimate the actual sensitivity of the values ($p_{\text{sens}}$) given their detected sensitive type ($t_{i}$) and internal context of the full document ($C_{I}$), to finally turn into a discrete sensitivity label ($s_{i}$). This step is crucial for mitigating false positives from the type detection stage by incorporating inter-value relationships and document semantics, to filter out values of sensitive types that are not actually sensitive.
Tabular Data case.
Figure 1 illustrates the detect-then-reflect mechanism for type contextualization on a concrete example with tabular data. Here, table columns are classified to correspond to any sensitive types such as address and phone number based on the column name and data values. For columns of sensitive types, the reflection step assesses whether the column is sensitive in context of the full table context, i.e., the table name as well as column names and values of surrounding columns.
@@ 3.1 Type Contextualization
Figure 1: Type contextualization for contextually detecting sensitive table columns using the detect-then-reflect mechanism to increase precision. This mechanism entails: ➀ detecting potentially sensitive columns based on their (semantic) type, ➁ reflecting on the actual sensitivity of the column within the internal context of the entire table.
3.2 Domain Contextualization
Besides a more fine-grained contextual notion of the sensitivity of certain values, we also argue for a more coarse-grained contextual perspective. Following our notion of contextual sensitive data (Def. 1), we pose that data protection should account for risks beyond identity exposure, for example, harming vulnerable populations or organizations hence consider the data domain, from humanitarian aid to national security. We refer to this broader perspective on data sensitivity as domain contextualization: sensitivity emerges from the dataset’s nature and potential misuse, which can arise from external factors. As such, domain contextualization aims at higher detection coverage by grounding the sensitivity assessment in domain-specific context, beyond generic and static predefined types.
The retrieve-then-detect mechanism
While type-based data sensitivity can often be inferred from internal document context, domain-specific sensitivity typically depends on contextual factors external to the data itself. For domain contextualization, we introduce the retrieve-then-detect mechanism (Alg. 2).
Algorithm 2 Domain Contextualization
1:Values $V_{i}=\{v_{1},\dots,v_{N}\}$ from document $D$
2:External context $C_{E}$, threshold $t$
[…]
Unlike type-informed detection, based on a predefined set of sensitive types, domain-specific sensitive data has no fixed ontology. Instead, the retrieve-then-detect mechanism considers external contextual information, such as rule-based heuristics specified in data governance documents or geopolitical news items, in the detection process. In the retrieval stage, external context is selected from a corpus of domain-specific documents or other external data sources ($C_{E}$) based on the relevance to the potential sensitive values at hand ($V_{i}$). In the detection stage, the sensitivity of the values at hand ($s_{i}$) is assessed given the retrieved relevant external context ($C_{i}^{\text{relevant}}$).
Tabular Data case.
Figure 2 illustrates the retrieve-then-detect mechanism for tabular data. For a given column (or table), the retrieval component fetches information, such as data policy rules or geopolitical information, from external documents or other data sources that are relevant for the given input table. The detection step then assesses whether a table (column), or combination of columns, contains data that should be considered sensitive given the retrieved relevant context.
@@ 3.2 Domain Contextualization
Figure 2: The retrieve-then-detect mechanism for detecting sensitive data that requires domain contextualization. This mechanism entails: ➀ retrieving relevant contextual information such as data policy rules, ➁ detecting sensitive data based on the table and the retrieved relevant context.